Find the angle between A=2i+2j-k and B=6i-3j+2k
1
Expert's answer
2019-10-21T10:50:10-0400

Let we have two planes given by vectors "A = A_1i+B_1j+C_1k" and "B = A_2i+B_2j+C_2k".

In our case

"A_1 = 2; B_1 = 2; C_1 = -1;"

"A_2 = 6; B_2 = -3; C_2 = 2"

The angle between the planes is equal to the angle between their normal vectors, it is enough to find the angle between the vectors "n_1(A_1,B_1,C_1)" and "n_2(A_2,B_2,C_2)".

That is

"\cos(\phi) = \cfrac{A_1A_2 + B_1B_2+C_1C_2 }{\sqrt{A_1^2+B_1^2+C_1^2}\sqrt{A_2^2+B_2^2+C_2^2}}"

"\cos(\phi) = \cfrac{2*6 + 2*(-3)+(-1)*2 }{\sqrt{2^2+2^2+(-1)^2}\sqrt{6^2+(-3)^2+2^2}} = \cfrac{4}{\sqrt{9}\sqrt{49}} = \cfrac{4}{21} \approx 0.19047619"

"\phi = \arccos(0.19047619) \approx 79.0194 \degree"

Answer: the angle between "A=2i+2j-k" and "B=6i-3j+2k" is "79.0194 \degree".